4.) A gas sample is held at constant pressure. The gas occupies 3.62 L of volume when the temperature is 21.6°C. Determine the temperature at which the volume of the gas is 3.38 L.
Answer #1

According to the ideal gas equation we have,

PV = nRT

So, at constant pressure;

V1/T1 = V2/T2

V1 = volume of gas at the temperature T1 (21.6°C = 294.6K)

V2 = volume of gas at the temperature T2

so,  3.62 L/ 294.6K = 3.38 L/ T2

T2 = 275.06K (= 2.06°C)

Hope this helps!